OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope intercept form: y + 2 = –2(x – 5)

OpenStudy (anonymous):

y+2 = -2(x-5) y+2 = -2x + 10 if u can solve from here. basically it should be of the form y = mx + c. ( or b, whatever u call it)

OpenStudy (anonymous):

i don't know how to solve it

OpenStudy (anonymous):

well the thing is, y+2 = -2x + 10 y = -2x + 8 . This is the slope-intercept form of the equation y = mx + c, where m = -2, c = 8.

OpenStudy (anonymous):

okay. where do you get the ten?

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

-2(x-5) = -2x + 10.